OpenStudy (anonymous):

percent of increase 2b to 5b

OpenStudy (anonymous):

visit this for better explanation- http://www.themathpage.com/arith/percent-increase-or-decrease.htm

hero (hero):

General formula for percent change \[\frac{|a - b|}{a} \times 100\] a = first number b = second number

OpenStudy (anonymous):

Yes 150% my bad.

hero (hero):

@Luis_Rivera , good to see your name working for the ping but there is nothing wrong with my formula. If you use my formula, you get the same answer. There's nothing to settle. I'm not trying to argue with you, but I think you posted that because you assumed what I posted was wrong.

hero (hero):

@Luis_Rivera, so you acknowledge that my formula is correct, right?